OpenStudy (anonymous):

IS If f(x) = x2 + x and g(x) = x - 3, find f(g(7)). IS THE ANWSER 28 ???? IF NOT CAN SOMEONE HELP ME

OpenStudy (anonymous):

WIO I GOT 28??

OpenStudy (anonymous):

is that right ??

OpenStudy (anonymous):

WHAT? YOU GOT 27?

OpenStudy (anonymous):

no i got 28 as the answer for If f(x) = x2 + x and g(x) = x - 3, find f(g(7)).

OpenStudy (anonymous):

Do you know what \(g(7)\)?

OpenStudy (anonymous):

is it 7

OpenStudy (anonymous):

Isn't \(g(x) = x-3\)?

OpenStudy (anonymous):

yes so x=7 right? so g(7)

OpenStudy (anonymous):

Then \(g(7) = 7-3\)

OpenStudy (anonymous):

CAN YOU SIMPLIFY?

OpenStudy (anonymous):

its 4 but the answer choices are 20,28,32 or53

OpenStudy (anonymous):

there is no 4

OpenStudy (anonymous):

The answer is not 28.

OpenStudy (anonymous):

So if you have \(g(7)\), then can you find \(f(g(7))\)?

OpenStudy (anonymous):

ya u can ?

OpenStudy (anonymous):

SO WHAT DO YOU GET?

OpenStudy (anonymous):

First plug x=7 into g(x)=x-3

OpenStudy (anonymous):

i got 28 when i plug the numbers in

OpenStudy (anonymous):

you should get G(7)=7-3

OpenStudy (anonymous):

SHOW YOUR WORK

OpenStudy (anonymous):

which equals 4

OpenStudy (anonymous):

yea but the answer choices are 20 28 32 53

OpenStudy (anonymous):

WHAT IS F(4)?

OpenStudy (anonymous):

so g(7) would equal 4 but were not done yet

OpenStudy (anonymous):

WHO SAID G(7) WAS THE ANSWER?

OpenStudy (anonymous):

now you plug g(7) into f(x) and you should get

OpenStudy (anonymous):

ohh ok so its 20???

OpenStudy (anonymous):

\[ f(g(7)) = f((7)-3) = f(4) = (4)^2+(4) \]

OpenStudy (anonymous):

f(g(7))= f(4)= 4^2+4=16=4 which would equal 20

OpenStudy (anonymous):

ohh so its 20 ohh so when u get 4 u plug that in ?? also

OpenStudy (anonymous):

SOOOO f(g(7)) equals 20

OpenStudy (anonymous):

yes plug in the 4

OpenStudy (anonymous):

G(7)=7-3. &-3 equals 4. Thats where you get the 4 and you plug it in.

OpenStudy (anonymous):

ohh so If f(x) = x2 and g(x) = x + 6, find g(f(0)) it will be =6 right because its 0??

OpenStudy (anonymous):

???

OpenStudy (anonymous):

yea.

OpenStudy (anonymous):

so the answer for If f(x) = x2 and g(x) = x + 6, find g(f(0)) is 6 because you plug n 0 for X ??????

OpenStudy (anonymous):

????

OpenStudy (anonymous):

That question is not the question you asked.

OpenStudy (anonymous):

post another thread for it.

OpenStudy (anonymous):

no i know I'm getting examples so i can learn how to do it so my example is If f(x) = x2 and g(x) = x + 6, find g(f(0)) i say the answer is 6